# Stabilizing Non-Abelian Topological Order against Heralded Noise via Local Lindbladian Dynamics

### v1

@@ VI Conclusions
Second, our local decoder exploits the fact that the fusion rules of the $D_{4}$ topological order are acyclic, meaning that two non-Abelian anyons of the same type fuse only to Abelian anyons. This property allows us to remove non-Abelian defects by applying a sequence of single-site $X$ operators. While advantageous for a general non-Abelian decoder (such as the MWPM protocol used in Sec. IV), this feature is essential for our local decoder, since the string operators that pair-annihilate non-Abelian anyons without generating further excitations along their length are high-depth unitary circuits that cannot be decomposed into a sequence of purely local operations applied independent of the correction history. In replacing these high-depth unitary circuits with strings of onsite $X$ operators, the main price that we pay is the creation of large numbers of Abelian anyons along the strings’ length; these however can be corrected locally at sufficiently high $Z$-correction rates. In principle, we expect that our correction protocol can be generalized to other acyclic anyon theories e.g., those described by $\mathcal{D}(G)$, where $G$ is some class-$n$ nilpotency group (where $n$ is finite). It would be interesting to investigate the conditions under which these generalizations also admit fully active phases that remain topologically ordered, leading to self-correcting quantum memories when the noise is perfectly heralded. It remains an open question whether our protocol generalizes to the case when $G$ is solvable but not nilpotent (e.g., $S_{3}$), where state preparation remains efficient under measurement and feedforward [82]; this question is particularly interesting since anyon theories based on solvable but not nilpotent groups $G$ can be used for universal topological quantum computation [91].
@@
One might also wonder about the feasibility of achieving the correction rates necessary to stabilize the TO phase on current hardware, since the relative rate of correction processes $\gamma/\eta\approx 500$ is very large. In practice, however, the measurements are performed only when the classical erasure flags are in an appropriate correctable configuration. Thus the rate at which measurement and feedback operations are performed, particularly for $Z$-correction operations, is far lower than the bare error correction rate. Fig. 17 shows the rate at which measurement and feedback operations relevant for $Z$ correction are performed (i.e. measurement of $A_{p}$) at various points in the phase diagram, revealing that the rate at which operations are actually applied is much lower than the bare correction rates throughout the active phase.
@@ Acknowledgements.
A.P. is grateful to Ramanjit Sohal and Nat Tantivasadakarn for helpful conversations. S.G. thanks Curt von Keyserlingk, Tibor Rakovszky, and Jeff Thompson for helpful discussions. F.J.B. and S.C. are grateful for the support of NSF DMR-2313858. S.G. acknowledges funding from an Institute for Robust Quantum Simulation (RQS) seed grant. This material is based upon work supported by the Sivian Fund and the Paul Dirac Fund at the Institute for Advanced Study and the U.S. Department of Energy, Office of Science, Office of High Energy Physics under Award Number DE-SC0009988 (A.P.). The authors acknowledge the computational resources provided by the Minnesota Supercomputing Institute (MSI) at the University of Minnesota.
Figure 18: The change in the total flag density $\delta[n^{f}]$ (see Eq. (45)) over a time interval $\Delta t$ is shown for (a) the Toric code (b) the $D_{4}$ model in the $\gamma_{x}-\gamma_{z}$ parameter space at some late time $t_{f}$. In both plots, the change in the value of density after time $\Delta t$ is small for parameters away from the phase boundaries. This suggest that the late time values evaluated at $t_{f}$ (see corresponding phase diagrams for flag densities in Fig. 1) are a good estimate of the steady-state densities. Near the transition, determination of the steady-state values require more careful analysis of larger system sizes (presented for one specific value $\gamma_{z}=500$ in Fig. 14 to identify the location of transition).
@@ Appendix A Details of numerical Monte Carlo simulations
         with probability $32c_{0}\gamma_{x}/$ 3: apply the appropriate jump operator from the $X$ correction move-set that is consistent with the flag structure at vertex $v_{j}$
@@ Appendix B Review of $D_{4}$ Topological Order
with center $\langle s^{2}\rangle$. There are five conjugacy classes $\{1\}$, $\{s^{2}\}$, $\{s,s^{3}\}$, $\{r,rs^{2}\}$, $\{rs,rs^{3}\}$, and hence also five irreducible representations (irreps): four of dimension one and one of dimension two. We label the one dimensional irreps by $ij$ ($i,j\in\mathbb{Z}_{2}$). Besides the trivial irrep $00$ and the faithful two dimensional irrep $\sigma$, there are three sign irreps:
@@
The anyon content of $D_{4}$ TO is specified as follows: each point-like anyon can be labelled by the ordered pair $([g],\pi_{g})$, where $[g]$ corresponds to a conjugacy class and $\pi_{g}$ denotes an irrep of its centralizer [90]. Thus, the four pure flux excitations are locally labelled by the non-trivial conjugacy classes, while the four pure charge excitations are labelled by the non-trivial elements of Rep$(D_{4})$. The remaining thirteen non-trivial excitations are dyons, leading to a total of twenty-two anyons (including the trivial anyon) in $D_{4}$ TO. The braiding phase between the pure charges and fluxes is given by the phase of the corresponding entry in the character table above: for instance, the wave function acquires a minus sign upon braiding the $11$ charge around the $[s]$ flux or when the $\sigma$ charge is braided with the $[s^{2}]$ flux.
@@
$1$ $D_{4}$ $00$ 1 1
$1$ $D_{4}$ $10$ $e_{rg}$ 1
$1$ $D_{4}$ $01$ $e_{r}$ 1
$1$ $D_{4}$ $11$ $e_{g}$ 1
$1$ $D_{4}$ $\sigma$ $m_{b}$ 2
$[{s^{2}}]$ $D_{4}$ $00$ $e_{rgb}$ 1
$[{s^{2}}]$ $D_{4}$ $10$ $e_{b}$ 1
$[{s^{2}}$] $D_{4}$ $01$ $e_{gb}$ 1
$[{s^{2}}]$ $D_{4}$ $11$ $e_{rb}$ 1
$[{s^{2}}]$ $D_{4}$ $\sigma$ $\psi_{b}$ 2
$[{s}]$ $\mathbb{Z}_{4}$ $1$ $m_{rg}$ 2
$[{s}]$ $\mathbb{Z}_{4}$ ${\mathbf{\omega}}$ $s_{rgb}$ 2
$[{s}]$ $\mathbb{Z}_{4}$ ${\mathbf{\omega}^{2}}$ $\psi_{rg}$ 2
$[{s}]$ $\mathbb{Z}_{4}$ $\overline{{\mathbf{\omega}}}$ $\bar{s}_{rgb}$ 2
$[{r}]$ $\mathbb{Z}_{2}^{2}$ $1$ $m_{gb}$ 2
$[{r}]$ $\mathbb{Z}_{2}^{2}$ $(-1,1)$ $m_{g}$ 2
$[{r}]$ $\mathbb{Z}_{2}^{2}$ $(1,-1)$ $\psi_{g}$ 2
$[{r}]$ $\mathbb{Z}_{2}^{2}$ $(-1,-1)$ $\psi_{gb}$ 2
$[{sr}]$ $\mathbb{Z}_{2}^{2}$ $1$ $m_{rb}$ 2
$[{sr}]$ $\mathbb{Z}_{2}^{2}$ $(-1,1)$ $m_{r}$ 2
$[{sr}]$ $\mathbb{Z}_{2}^{2}$ $(1,-1)$ $\psi_{r}$ 2
$[{sr}]$ $\mathbb{Z}_{2}^{2}$ $(-1,-1)$ $\psi_{rb}$ 2
@@ Appendix C Simulation using Non-Pauli stabilizer tableau
In this Appendix, we provide details for simulating the Lindbladian dynamics of the $D_{4}$ model defined in terms of non-Pauli stabilizers (see Eq. (20). For an $N$ qubit system, the time evolution of standard Pauli stabilizers (and consequently, the states that they stabilize) can be simulated by tracking the time evolution of the $O(N)$ stabilizers, each of which is a local product of Pauli operators. Since dynamics generated by Clifford gates maps Pauli stabilizers to Pauli stabilizers, any Clifford-based dynamics can be simulated efficiently by simply updating the stabilizers at each time step [92]. The ground-space of the $D_{4}$ model is instead stabilized by the non-Pauli stabilizers (or “quasi-stabilizers”) defined in Eqs. (18) and (19): each of these is a tensor-product of $1$ and $2$ qubit operators that are localized around a single vertex or plaquette of the honeycomb lattice. In an abuse of terminology, we will refer to these operators simply as stabilizers but we stress that these are not mutually commuting elements of the $N$-qubit Pauli group. We now provide a generalized protocol that shows how the evolution of these stabilizers under the Pauli operations present in our dynamical protocol can be tracked similarly to genuine stabilizers, allowing us to efficiently simulate the noise and correction Lindbladian dynamics of the $D_{4}$ model.
We begin by choosing a basis for the independent stabilizers of the initial $D_{4}$ state. Any ground state of the $D_{4}$ Hamiltonian Eq. (20) is fully fixed by a list of plaquette and vertex stabilizer generators, together with the non-local logical operators that stabilize it. Each vertex of the honeycomb lattice has a vertex stabilizer $B_{v}$ (see Eq. (18)). While there are $32L^{2}/$ 3 operators on each sublattice of a system with $L\times L$ lattice sites, for each color $c$ these obey a global constraint $\prod_{v_{c}}B_{v_{c}}=\mathbb{I}$. Hence, the set of independent generators can be represented by
@@
$\mathfrak{L}=\begin{cases}\{\mathcal{Z}^{r}_{V},\mathcal{Z}^{g}_{V},\mathcal{Z}^{b}_{V},\mathcal{Z}^{r}_{H},\mathcal{Z}^{g}_{H},\mathcal{Z}^{b}_{H}\},\ \text{ if }|\psi_{0}\rangle=|\widetilde{0}\rangle_{L}\\ \{\mathcal{Z}^{r}_{V},\mathcal{Z}^{g}_{V},\mathcal{Z}^{b}_{V},\mathcal{X}^{r}_{V},\mathcal{X}^{g}_{V},\mathcal{X}^{b}_{V}\},\ \text{ if }|\psi_{0}\rangle=|\widetilde{+}\rangle_{L}\\ \end{cases}$ (58)
In total, the $3L^{2}$ operators from the combined sets $\mathcal{Q},\mathcal{G}$, and $\mathfrak{L}$ describe a unique state as their $+1$ eigenvector [64].
@@ C.1 Evolution under $X$ and $Z$ operators
To simulate the dynamics, we will track individual trajectories of our system over time. If $S$ stabilizes the state $|\psi(t)\rangle$ at time $t$, then the state $|\psi(t+1)\rangle:=U|\psi(t)\rangle$ is stabilized by the modified stabilizer operator $S^{\prime}:=USU^{-1}$. This unitary evolution also affects the destabilizer operators, which also need to be updated according to $D^{\prime}:=UDU^{-1}$. This simultaneous transformation leaves the defining properties of destabilizers in Eq. (59) invariant. In the case at hand, both the noise and feedback operations involve application of Pauli $X$ and $Z$ operators. This leads to the modified $B_{v}$ and $A_{p}$ operators given in Eq. (36), (37), and (38).
@@
±$\pm\mathcal{Z}\xrightarrow{X_{k}}\mp\mathcal{Z}$ (61)
@@
±$\pm\mathcal{X}\xrightarrow{Z_{k}}\mp\mathcal{X},\text{ and}\ \pm\mathcal{X}\xrightarrow{X_{j}}\mp\prod_{i\in\xi(j)}Z_{i}\mathcal{X}.$ (62)
@@ C.2 Measurement of $B_{v}$ and $A_{p}$
The vertex operators $B_{v}$ commute with all stabilizers and hence their measurement leaves the state unchanged. The corresponding measurement outcome is readily available from the sign of elements in $\mathcal{Q}$. Similarly, if a measurement operator $A_{p}$ is present in the list of instantaneous generators (i.e. ±$\pm A_{p}\in\mathcal{G}$), then the outcome is given by the sign of the stabilizer and the state is not modified.
If ±$\pm A_{p}$ is not amongst the list of generators, then we need to observe its algebraic relation with the generators in order to determine the result of the measurement on the state. In particular, note that $A_{p^{\prime}}$ commutes with $A_{p}$ up to a product of vertex operators (see eg. Eq. (22)). It follows that the instantaneous stabilizer generators $g=\prod_{k}Z_{k}\prod_{p^{\prime}}A_{p^{\prime}}$ (that is potentially made up of multiple $A$ type operators) obey the relation:
$\begin{split}A_{p}g&=A_{p}(\prod_{k}Z_{k}\prod_{p^{\prime}}A_{p^{\prime}})=(\prod_{k}Z_{k}A_{p}^{\prime})A_{p}(\mathsf{s}\prod_{v\in\mathcal{V}}B_{v})\\ &=gA_{p}\ (\mathsf{s}\prod_{v\in\mathcal{V}}B_{v})\,,\end{split}$ (64)
@@
where in the last equality, we have used the fact that $\mathsf{p_{-m}}:=\langle\psi|(1-\mathsf{m}A_{p})/2|\psi\rangle$. Since $A_{p}$ is a hermitian operator that squares to $1$, $\mathsf{m}=\pm 1$ are the only possible measurement outcomes which, together with Eq. (66), implies that $\mathsf{p}_{+1}=\mathsf{p}_{-1}=1/2$. Hence, the measurement outcome in this case is equally likely to be $+1$ or $-1$, which we simulate by sampling a binary random number from the uniform distribution. Suppose $g_{1},g_{2},\ldots,g_{n_{0}}$ are the $n_{0}\geq 1$ stabilizers for which this holds. Then, they no longer stabilize the post-measurement state stabilized by $\mathsf{m}A_{p}$ since they obey $g_{j}(1+\mathsf{m}A_{p})|\psi\rangle=(1-\mathsf{m}A_{p})|\psi\rangle$. Instead, the new state of the system is stabilized by pairwise products of these generators. This leads to a modification of stabilizers as follows: we choose a particular stabilizer $g_{1}$ from this list and the remaining anti-commuting operators are updated as
@@
where the variables $\mu,\nu,\xi\in\{0,1\}$ denote the presence or absence of the corresponding generator in the operator string. Recall from the definitions in Eqs. (55) and (57) that the individual operators $q$ and $\mathcal{Z}$ are products of Pauli-Z operators whereas the operators $g$ additionally involve plaquette stabilizers $A$. The goal is to find out the value of $m$ by determining which generators multiply together to give $mA_{p}$.
@@
$\begin{split}\mathsf{m}=&\big{(}A_{p}\prod_{i:\mu_{i}=1}g_{i}\big{)}\ \prod_{j}(\mathsf{s_{q_{j}}}B_{j})^{\nu_{j}}\\ &=\big{(}\mathsf{s^{\prime}}\prod_{l\in\mathcal{E}}Z_{l}\prod_{j}\big{)}\ (\mathsf{s_{q_{j}}}B_{j})^{\nu_{j}}=\mathsf{s^{\prime}}\times\prod_{j:\nu_{j}=1}\mathsf{s_{q_{j}}}\end{split}$ (71)
@@
The update of logical operators as a result of the measurements are carried out similarly. The $Z$ logical operators commute with both $B_{v}$ and $A_{p}$ operators and hence remain unchanged. The logical $X$ operators have non-trivial commutation relations with $A_{p}$ as a result of Eq. (24). If, at a certain time step, measurement operators $A_{p}$ anticommute with the logical (i.e. $\{A_{p},\mathcal{X}_{L}\}|\psi\rangle=0$) in addition to some other stabilizers $g$, then we update the logical $X$ operator as discussed in case a) above. If we find that the measurement only anti-commutes with the logical $X$ stabilizer, then it would imply that the state is no longer stabilized by the logical $X$ operator after the measurement. We term it as a logical error since it collapses the logical $X$ eigenstate.

### v2

@@ VI Conclusions
Second, our local decoder exploits the fact that the fusion rules of the $D_{4}$ topological order are acyclic, meaning that two non-Abelian anyons of the same type fuse only to Abelian anyons. This property allows us to remove non-Abelian defects by applying a sequence of single-site $X$ operators. While advantageous for a general non-Abelian decoder (such as the MWPM protocol used in Sec. IV), this feature is essential for our local decoder, since the string operators that pair-annihilate non-Abelian anyons without generating further excitations along their length are high-depth unitary circuits that cannot be decomposed into a sequence of purely local operations applied independent of the correction history. In replacing these high-depth unitary circuits with strings of onsite $X$ operators, the main price that we pay is the creation of large numbers of Abelian anyons along the strings’ length; these however can be corrected locally at sufficiently high $Z$-correction rates. In principle, we expect that our correction protocol can be generalized to other acyclic anyon theories e.g., those described by $\mathcal{D}(G)$, where $G$ is some class-$n$ nilpotency group (where $n$ is finite). It would be interesting to investigate the conditions under which these generalizations also admit fully active phases that remain topologically ordered, leading to self-correcting quantum memories when the noise is perfectly heralded. It remains an open question whether our protocol generalizes to the case when $G$ is solvable but not nilpotent (e.g., $S_{3}$), where state preparation remains efficient under measurement and feedforward [86]; this question is particularly interesting since anyon theories based on solvable but not nilpotent groups $G$ can be used for universal topological quantum computation [113, 95].
@@
Our results also provide insight into steady-state phases for models in a TO phase away from their fixed points (such as the Toric code or $D_{4}$ model discussed here). The ground state of any point in the TO phase is connected to the respective fixed-point ground state by finite-depth quasi-local unitary operators [114, 115, 1]. This, in turn, means that the code-space of such a general point in the TO phase will be defined by quasi-local quasi-stabilizers. As a result, single-qubit errors will create more quasi-stabilizer violations compared to the fixed-point state. However, if these errors are perfectly heralded, then the resulting quasi-stabilizer violations will still be connected by strings of flags on some coarse-grained level. We expect that an appropriately modified version of our protocol can successfully target a stable active phase even in this case.
From a practical point of view, we also briefly commented on the feasibility of achieving the correction rates necessary to stabilize the TO phase on current hardware. It turns out that the quantum memory phase requires a relatively large rate $\gamma/\eta\approx 500$ of correction processes involving entangling gates. In practice, however, the stabilizer measurements in our protocol have to be performed only when the classical erasure flags are in an appropriate correctable configuration. Thus, the actual rate at which the entangling gates are implemented, particularly during $Z$-correction protocol, is expected to be lower than the bare correction rate $\gamma_{z}$. In Fig. 17 we show the rate at which the measurement and feedback operations relevant for $Z$ correction are performed (i.e., measurement of $A_{p}$) at various points in the phase diagram, revealing that the actual gate implementation rates are much lower than the bare correction rates throughout the active phase.
@@ Acknowledgements.
A.P. is grateful to Ramanjit Sohal and Nat Tantivasadakarn for helpful conversations. S.G. thanks Curt von Keyserlingk, Tibor Rakovszky, and Jeff Thompson for helpful discussions. The authors acknowledge helpful conversations with David Huse. F.J.B. and S.C. are grateful for the support of NSF DMR-2313858. S.G. acknowledges funding from an Institute for Robust Quantum Simulation (RQS) seed grant. This material is based upon work supported by the Sivian Fund and the Paul Dirac Fund at the Institute for Advanced Study and the U.S. Department of Energy, Office of Science, Office of High Energy Physics under Award Number DE-SC0009988 (A.P.). The authors acknowledge the computational resources provided by the Minnesota Supercomputing Institute (MSI) at the University of Minnesota.
Data availability
All simulation codes and generated datasets are available online at [116].
Figure 18: The change in the total flag density $\delta[n^{f}]$ (see Eq. (46)) over a time interval $\Delta t$ is shown for (a) the Toric code (b) the $D_{4}$ model in the $\gamma_{x}-\gamma_{z}$ parameter space at some late time $t_{f}$. In both plots, the change in the value of density after time $\Delta t$ is small for parameters away from the phase boundaries. This suggests that the late time values evaluated at $t_{f}$ (see corresponding phase diagrams for flag densities in Fig. 1) are a good estimate of the steady-state densities. Near the transition, determination of the steady-state values requires more careful analysis of larger system sizes (presented for one specific value $\gamma_{z}=500$ in Fig. 14 to identify the location of transition).
@@ Appendix A Details of numerical Monte Carlo simulations
   with probability $2c_{0}\gamma_{x}/3$: apply the appropriate jump operator from the $X$ correction move-set that is consistent with the flag structure at vertex $v_{j}$
@@ Appendix B Review of $D_{4}$ Topological Order
with center $\langle s^{2}\rangle$. There are five conjugacy classes $\{1\}$, $\{s^{2}\}$, $\{s,s^{3}\}$, $\{r,rs^{2}\}$, $\{rs,rs^{3}\}$, and hence also five irreducible representations (irreps): four of dimension one and one of dimension two. We label the one dimensional irreps by ijij ($i,j\in\mathbb{Z}_{2}$). Besides the trivial irrep $00$ and the faithful two dimensional irrep $\sigma$, there are three sign irreps:
@@ 11 $^{2} ss$ $r$ rsrs
[…]
@@
11 $D_{4}$ $00$ 1 1
11 $D_{4}$ $10$ $e_{rg}$ 1
11 $D_{4}$ $01$ $e_{r}$ 1
11 $D_{4}$ $11$ $e_{g}$ 1
11 $D_{4}$ $\sigma$ $m_{b}$ 2
$s^{2}$}] $D_{4}$ $00$ $e_{rgb}$ 1
$s^{2}$}] $D_{4}$ $10$ $e_{b}$ 1
$s^{2}$}] $D_{4}$ $01$ $e_{gb}$ 1
$s^{2}$}] $D_{4}$ $11$ $e_{rb}$ 1
$s^{2}$}] $D_{4}$ $\sigma$ $\psi_{b}$ 2
[s][{s}] $\mathbb{Z}_{4}$ 11 $m_{rg}$ 2
[s][{s}] $\mathbb{Z}_{4}$ $\mathbf{\omega}$} $s_{rgb}$ 2
[s][{s}] $\mathbb{Z}_{4}$ $\mathbf{\omega}^{2}$} $\psi_{rg}$ 2
[s][{s}] $\mathbb{Z}_{4}$ $\overline{{\mathbf{\omega}}}$ $\bar{s}_{rgb}$ 2
[r][{r}] $\mathbb{Z}_{2}^{2}$ 11 $m_{gb}$ 2
[r][{r}] $\mathbb{Z}_{2}^{2}$ (1,1)(-1,1) $m_{g}$ 2
[r][{r}] $\mathbb{Z}_{2}^{2}$ (1,1)(1,-1) $\psi_{g}$ 2
[r][{r}] $\mathbb{Z}_{2}^{2}$ (1,1)(-1,-1) $\psi_{gb}$ 2
[sr][{sr}] $\mathbb{Z}_{2}^{2}$ 11 $m_{rb}$ 2
[sr][{sr}] $\mathbb{Z}_{2}^{2}$ (1,1)(-1,1) $m_{r}$ 2
[sr][{sr}] $\mathbb{Z}_{2}^{2}$ (1,1)(1,-1) $\psi_{r}$ 2
[sr][{sr}] $\mathbb{Z}_{2}^{2}$ (1,1)(-1,-1) $\psi_{rb}$ 2
@@ Appendix C Simulation using Non-Pauli stabilizer tableau
In this Appendix, we provide details for simulating the Lindbladian dynamics of the $D_{4}$ model defined in terms of non-Pauli stabilizers (see Eq. (21)). For an $N$ qubit system, the time evolution of standard Pauli stabilizers (and consequently, the states that they stabilize) can be simulated by tracking the time evolution of the $O(N)$ stabilizers, each of which is a local product of Pauli operators. Since dynamics generated by Clifford gates maps Pauli stabilizers to Pauli stabilizers, any Clifford-based dynamics can be simulated efficiently by simply updating the stabilizers at each time step [96]. The ground-space of the $D_{4}$ model is instead stabilized by the non-Pauli stabilizers (or “quasi-stabilizers”) defined in Eqs. (19) and (20): each of these is a tensor-product of $1$ and $2$ qubit operators that are localized around a single vertex or plaquette of the honeycomb lattice. In an abuse of terminology, we will refer to these operators simply as stabilizers, but we stress that these are not mutually commuting elements of the $N$-qubit Pauli group. We now provide a generalized protocol that shows how the evolution of these stabilizers under the Pauli operations present in our dynamical protocol can be tracked similarly to genuine stabilizers, allowing us to efficiently simulate the noise and correction Lindbladian dynamics of the $D_{4}$ model.
We begin by choosing a basis for the independent stabilizers of the initial $D_{4}$ state. Any ground state of the $D_{4}$ Hamiltonian Eq. (21) is fully fixed by a list of plaquette and vertex stabilizer generators, together with the non-local logical operators that stabilize it. Each vertex of the honeycomb lattice has a vertex stabilizer $B_{v}$ (see Eq. (19)). While there are $2L^{2}/3$ operators on each sublattice of a system with $L\times L$ lattice sites, for each color $c$, these obey a global constraint $\prod_{v_{c}}B_{v_{c}}=\mathbb{I}$. Hence, the set of independent generators can be represented by
@@
[…]
In total, the $3L^{2}$ operators from the combined sets $\mathcal{Q},\mathcal{G}$, and $\mathfrak{L}$ describe a unique state as their +1+1 eigenvector [67].
@@ C.1 Evolution under $X$ and $Z$ operators
To simulate the dynamics, we will track individual trajectories of our system over time. If $S$ stabilizes the state $|\psi(t)\rangle$ at time $t$, then the state $|\psi(t+1)\rangle:=U|\psi(t)\rangle$ is stabilized by the modified stabilizer operator $S^{\prime}:=USU^{-1}$. This unitary evolution also affects the destabilizer operators, which also need to be updated according to $D^{\prime}:=UDU^{-1}$. This simultaneous transformation leaves the defining properties of destabilizers in Eq. (60) invariant. In the case at hand, both the noise and feedback operations involve the application of Pauli $X$ and $Z$ operators. This leads to the modified $B_{v}$ and $A_{p}$ operators given in Eq. (37), (38), and (39).
@@
$\pm\mathcal{Z}\xrightarrow{X_{k}}\mp\mathcal{Z}$ (62)
@@
$\pm\mathcal{X}\xrightarrow{Z_{k}}\mp\mathcal{X},\text{ and}\ \pm\mathcal{X}\xrightarrow{X_{j}}\mp\prod_{i\in\xi(j)}Z_{i}\mathcal{X}.$ (63)
@@ C.2 Measurement of $B_{v}$ and $A_{p}$
The vertex operators $B_{v}$ commute with all stabilizers and hence their measurement leaves the state unchanged. The corresponding measurement outcome is readily available from the sign of elements in $\mathcal{Q}$. Similarly, if a measurement operator $A_{p}$ is present in the list of instantaneous generators (i.e., $\pm A_{p}\in\mathcal{G}$), then the outcome is given by the sign of the stabilizer and the state is not modified.
If $\pm A_{p}$ is not amongst the list of generators, then we need to observe its algebraic relation with the generators in order to determine the result of the measurement on the state. In particular, note that $A_{p^{\prime}}$ commutes with $A_{p}$ up to a product of vertex operators (see eg. Eq. (23)). It follows that the instantaneous stabilizer generators $g=\prod_{k}Z_{k}\prod_{p^{\prime}}A_{p^{\prime}}$ (that is potentially made up of multiple $A$ type operators) obey the relation:
[…]
@@
where in the last equality, we have used the fact that $\mathsf{p_{-m}}:=\langle\psi|(1-\mathsf{m}A_{p})/2|\psi\rangle$. Since $A_{p}$ is a hermitian operator that squares to $1$, $\mathsf{m}=\pm 1$ are the only possible measurement outcomes which, together with Eq. (67), implies that $\mathsf{p}_{+1}=\mathsf{p}_{-1}=1/2$. Hence, the measurement outcome in this case is equally likely to be +1+1 or 1-1, which we simulate by sampling a binary random number from the uniform distribution. Suppose $g_{1},g_{2},\ldots,g_{n_{0}}$ are the $n_{0}\geq 1$ stabilizers for which this holds. Then, they no longer stabilize the post-measurement state stabilized by $\mathsf{m}A_{p}$ since they obey $g_{j}(1+\mathsf{m}A_{p})|\psi\rangle=(1-\mathsf{m}A_{p})|\psi\rangle$. Instead, the new state of the system is stabilized by pairwise products of these generators. This leads to a modification of stabilizers as follows: we choose a particular stabilizer $g_{1}$ from this list and the remaining anti-commuting operators are updated as
@@
where the variables $\mu,\nu,\xi\in\{0,1\}$ denote the presence or absence of the corresponding generator in the operator string. Recall from the definitions in Eqs. (56) and (58) that the individual operators $q$ and $\mathcal{Z}$ are products of Pauli-Z operators whereas the operators $g$ additionally involve plaquette stabilizers $A$. The goal is to find out the value of $\mathsf{m}$ by determining which generators multiply together to give $\mathsf{m}A_{p}$.
@@
$\begin{split}\mathsf{m}=&\big(A_{p}\prod_{i:\mu_{i}=1}g_{i}\big)\ \prod_{j}(\mathsf{s_{q_{j}}}B_{j})^{\nu_{j}}\\ &=\big(\mathsf{s^{\prime}}\prod_{l\in\mathcal{E}}Z_{l}\big)\prod_{j}\ (\mathsf{s_{q_{j}}}B_{j})^{\nu_{j}}=\mathsf{s^{\prime}}\times\prod_{j:\nu_{j}=1}\mathsf{s_{q_{j}}}\end{split}$ (72)
@@
The updates of logical operators as a result of the measurements are carried out similarly. The $Z$ logical operators commute with both $B_{v}$ and $A_{p}$ operators and hence remain unchanged. The logical $X$ operators have non-trivial commutation relations with $A_{p}$ as a result of Eq. (25). If, at a certain time step, measurement operators $A_{p}$ anticommute with the logical (i.e. $\{A_{p},\mathcal{X}_{L}\}|\psi\rangle=0$) in addition to some other stabilizers $g$, then we update the logical $X$ operator as discussed in case a) above. If we find that the measurement only anti-commutes with the logical $X$ stabilizer, then it would imply that the state is no longer stabilized by the logical $X$ operator after the measurement. We term it as a logical error since it collapses the logical $X$ eigenstate.